40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 19 18:41:42 2022, yuta, Update, BHD, Contrast measurements for Michelson and ITM-LO ContrastMeasurements.pdf
    Reply  Wed Jul 20 18:07:52 2022, Paco, Update, BHD, BHD MICH test BHD_MICH_OSC.pdf
       Reply  Tue Jul 26 20:54:08 2022, Paco, Update, BHD, BHD MICH test - LO phase control 20220726_BICHD.pdf
          Reply  Wed Jul 27 18:30:50 2022, yuta, Update, BHD, LO beam power at BHD DCPDs is significantly lower than expected LOBeamAtBHD.JPGITMXSingleBounceAtBHD.jpg
             Reply  Thu Jul 28 14:34:40 2022, Yehonathan, Update, BHD, LO beam power at BHD DCPDs is significantly lower than expected 
                Reply  Thu Jul 28 16:51:55 2022, Tega, Update, BHD, Shaking test for LO beam AS beam to BHD DCPDs Screenshot_2022-07-28_16-58-34_LOandASShaking.pngScreenshot_2022-07-28_18-08-55_DCPDPOPSuspensionCoherence.png
                   Reply  Thu Jul 28 20:16:26 2022, Anchal, Update, BHD, Shaking test for LO beam AS beam to BHD DCPDs 
                   Reply  Fri Jul 29 18:24:53 2022, Tega, Update, BHD, LO beam power improved by factor of 6 after LO and AS beam alignment LO_and_oldAS_settings.pngBHD_fringe_settings.pngScreenshot_2022-07-29_19-37-39.pngFromTheLeft-AS-POP-LO.JPG
                      Reply  Wed Aug 3 16:00:51 2022, yuta, Update, BHD, BHD fringe aligned with reduced LO and AS beam clipping Screenshot_2022-08-03_15-46-36_BHDfringeAlmostUnclipped.png
                         Reply  Tue Aug 9 15:33:12 2022, yuta, Update, BHD, BHD fringe contrast improved from 43% to 74% HPC-DCPD_B_OUT_1344118517_ITMY-LO.pngHPC-DCPD_A_OUT_1344118517_ITMY-LO.pngHPC-DCPD_B_OUT_1344118318_ITMX-LO.pngHPC-DCPD_A_OUT_1344118318_ITMX-LO.png
                            Reply  Tue Aug 9 15:50:22 2022, Koji, Update, BHD, BHD fringe contrast improved from 43% to 74% 
                               Reply  Wed Nov 16 15:09:08 2022, yuta, Update, BHD, BHD fringe contrast measured with unwhitening filters ContrastMeasurements20221116_edited.pdf
                         Reply  Tue Oct 25 08:25:00 2022, JC, Update, BHD, BHD fringe aligned with reduced LO and AS beam clipping 
                         Reply  Thu Nov 17 11:24:39 2022, JC, HowTo, LSC, Locking MICH 
                            Reply  Tue Nov 29 11:38:37 2022, JC, HowTo, LSC, Lock Single Arm After MICH lock Screen_Shot_2022-11-29_at_11.01.56_AM.png
Message ID: 17046     Entry time: Fri Jul 29 18:24:53 2022     In reply to: 17044     Reply to this: 17056
Author: Tega 
Type: Update 
Category: BHD 
Subject: LO beam power improved by factor of 6 after LO and AS beam alignment 

[Yuta, Tega]

From our previous work (elog 17044) of shaking PR2 and seeing a signal in DCPD_A and the fact that LO beam power is far smaller than the expected nominal value, we decided to use TT1 and TT2 to realign the LO beam. This resulted in LO beam power going up by a factor of 6 and an improvement in the LO beam shape. We are still unable to find LO and AS alignment which realize BHD fringe with no clipping everywhere.

Deformed LO beam issue: Following the TT1 and TT2 alignments, used PR2 and PR3 to recover the transmission of the X and Y arms to 1. We also used LO1 and LO2 offsets to further reduce the beam deformation by eliminating the HOM concentric fringes that surounded the LO beam and to maximize the DCPD outputs. BHD optics in ITMY table was tweaked a lot to keep the LO beam centered on the BHD DCPDs and camera. The improved LO beam is still astigmatic in the yaw direction but at least now looks like a TEM00 mode. We also repositioned the DCPD_A path camera lens to remove the circular diffused fringes due to lens clipping. After the alignment, power was measured to be the following. It also reduced the coherence between DCPD outputs and suspension motions (see attached).

                       LO         ITMX
C1:HPC-DCPD_A_OUT16    +127.50    +96.24 (ITMX single bounce consistent to 40m/17040)
C1:HPC-DCPD_B_OUT16    +120.51    +141.52
Power at viewport A    504 uW (almost as expected 40m/17040)
Power at viewport B    385 uW

AP table AS beam clipping: We also noticed clipping in the AS beam in AP table which we removed by moving SR2 and AS1 in YAW and then used AS4 to keep the BHD AS beam centered in the BHD DCPDs.

BHD fringe: After overlaping the LO and AS beams, we saw diagonal fringes indicating beam tilt of LO wrt AS, so we tried to remove the AS beam tilt using AS1 and AS4 but failed to do so because the AS4 mirror seemed to completely distort the beam, so intead we decided to use SR2 and AS1 to remove the tilt between LO and AS beams, which realized BHD fringe. But the motion of SR2 and AS1 then moved the AS beam that it is no longer seen in AP table. The alignment to realize LO and AP AS beam without clipping, and that to realize BHD fringe are attached.

Attachment 1: LO_and_oldAS_settings.png  32 kB  Uploaded Fri Jul 29 20:15:04 2022  | Hide | Hide all | Show all
LO_and_oldAS_settings.png
Attachment 2: BHD_fringe_settings.png  30 kB  Uploaded Fri Jul 29 20:15:04 2022  | Show | Hide all | Show all
Attachment 3: Screenshot_2022-07-29_19-37-39.png  181 kB  Uploaded Fri Jul 29 20:38:35 2022  | Hide | Hide all | Show all
Screenshot_2022-07-29_19-37-39.png
Attachment 4: FromTheLeft-AS-POP-LO.JPG  3.012 MB  Uploaded Fri Jul 29 20:46:12 2022  | Hide | Hide all | Show all
FromTheLeft-AS-POP-LO.JPG
ELOG V3.1.3-